# Linear Simple Cycle Reservoirs at the edge of stability perform Fourier decomposition of the input driving signals

### v1

@@ 4. Unit spectral radius SCR implies harmonic motifs in Complex Domain
Here, by the assumption that $\tau=n$, we can define two $n\times n$ matrices $\mbox{\bf{D}}=\operatorname{diag}\{d_{1},d_{2},\cdots,d_{n}\}$ and $\mbox{\bf{F}}=[\omega_{i}^{j-1}]$. Note that F is the discrete Fourier transform matrix, so we know $\mbox{\bf{F}}^{*}\mbox{\bf{F}}=\mbox{\bf{F}}\mbox{\bf{F}}^{*}=n\mbox{\bf{I}}$. Therefore,
@@
When $\tau\neq n$, we can still reach a similar decomposition as above, but the matrix $\mbox{\bf{F}}=[\omega_{i}^{j-1}]$ is an $n\times\tau$ matrix. When $\tau$ is not an integer multiple of $n$ in particular, its rows may not be orthogonal to each other in general, so it is hard to characterize the eigenvectors of Q. However, when $\tau$ is an integer multiple of $n$, the rows of F are orthogonal to each other. In this case, the eigenvectors of Q is given by the $n$ of these $\tau$ dimensional vectors $(\omega_{i}^{j})_{j=0}^{\tau-1}$.

### v2

@@ Abstract.
Recurrent Neural Networks (RNNs) are machine learning methods for modeling temporal dependencies in sequential data, but their training can be computationally demanding. Reservoir Computing (RC), a simplified subset of RNNs, circumvents this problem by fixing the internal dynamics of the network (the reservoir) and focusing training on the readout layer. Simple Cycle Reservoir (SCR) is a type RC model that stands out for its minimalistic design and proven capability to universally approximate a wide class of processes operating on time series data (namely time-invariant fading memory filters) even in the linear dynamics regime (and non-linear static readouts). We show that, interestingly, when linear SCR is constructed at the edge of stability, it implicitly represents the time series according to a well known and widely used technique of Fourier signal decomposition. This insight demonstrates that deep connections can exist between recurrent neural networks, classical signal processing techniques and statistics, paving the way for their enhanced understanding and innovative applications.
@@ 4. Unit spectral radius SCR implies harmonic motifs in Complex Domain
Here, by the assumption that $\tau=n$, we can define two $n\times n$ matrices $\mbox{\bf{D}}=\operatorname{diag}\{d_{1},d_{2},\cdots,d_{n}\}$ and $\mbox{\bf{F}}=[\omega_{i}^{-(j-1)}]$. Note that $\mbox{\bf{F}}^{*}$ is the discrete Fourier transform matrix, so we know $\mbox{\bf{F}}^{*}\mbox{\bf{F}}=\mbox{\bf{F}}\mbox{\bf{F}}^{*}=n\mbox{\bf{I}}$. Therefore,
@@
When $\tau\neq n$, we can still reach a similar decomposition as above, but the matrix $\mbox{\bf{F}}=[\omega_{i}^{-(j-1)}]$ is an $n\times\tau$ matrix. When $\tau$ is not an integer multiple of $n$ in particular, its rows may not be orthogonal to each other in general, so it is hard to characterize the eigenvectors of Q. However, when $\tau$ is an integer multiple of $n$, the rows of F are orthogonal to each other. In this case, the eigenvectors of Q is given by the $n$ of these $\tau$ dimensional vectors $(\omega_{i}^{-j})_{j=0}^{\tau-1}$.